### v1

@@ Appendix C Activity Analysis
| GP of the SOPHIE optical RVs using SPIRou’s as activity proxy | ||||
@@
Appendix E Corner plots and additional figures and tables of the one-planet model.
Appendix F Corner plots and additional figures and tables of the RadVel 2-planet model.
@@
Appendix G Dynamical stability analysis
While somehow limited constraints on the orbital configuration can be obtained from the radial velocity measurements, we performed a dynamical stability assessment for the two-planet system Gl 410b and 18.7 day candidate to consolidate the picture on the system architecture. The period ratio (larger than 3) is wide and Gl 410 b and the 18.7d day candidate are small enough that the system is safely stable for circular orbits and the nominal eccentricity values. Indeed the orbital separation in Hill radii is of which is far beyond the stability limit of (Gladman 1993). Although we favor a model with circular orbits, we can derive some constraints on the maximum stable eccentricities by computing the general Hill stability criterion as described in Petit et al. (2018). We take the RadVel posterior and compute the Angular Momentum Deficit (AMD) of the system as well as the critical value ensuring Hill stability (in other words, the maximum AMD such that close encounters are always forbidden). In Fig. 30, we compare the posterior eccentricity distribution resulting from the RadVel RV fit and the posterior eccentricity distribution only keeping the stable configurations. Rejecting Hill unstable configurations from the posterior, we obtain an upper limit for at 0.4 but no additional constraints on . Stable configurations are on average more circular. A similar analysis on the masses show that no stringer constraints can be drawn from the stability on this system.
The Hill stability can also constrain the maximum mutual inclination. To obtain this constraint, we assume that the system invariant plan is aligned with the line of sight191919As the main constraint comes from the mutual inclination and not the impact on the real planetary masses, this is a reasonable hypothesis. and we assign a random mutual inclination to each configuration in the posterior. Computing the critical AMD in this case, we find that systems with a mutual inclination smaller than 40 degrees are stable. As for the eccentricities, this is mainly an upper limit. The system is fully consistent with coplanar orbits.
Appendix H Additional RV analysis of the SPIRou wPCA data.
H.1 periodogram analysis
The periodogram has several inputs: the RV time series to be analysed, its nominal uncertainties, a grid of frequencies , a linear model assumed to be in the data , an assumed covariance matrix of the noise as input. Overall, the model of the RV time series can be described by the equation
@@ Appendix G Dynamical stability analysis
The periodogram looks for a small number of frequencies such that the model in Eq. (3) approximates the data. Depending on the different assumptions on the noise model and base linear model, the periodogram might find signals at different frequencies. To test the robustness of the results to a change in the model, we adopt the methodology of Hara et al. (2020). We consider a grid of potential noise models. For each of them, we compute the periodogram, and keep the signals that have a false alarm probability below 0.1. For each noise model in the grid, we then have candidate periodic signals. To each couple noise model/periodic signals, we attribute a cross-validation score. The model is fitted on 60% of the data points (the training set), and the likelihood of its prediction on the remaining 40% (the test set) is computed. This operation is repeated for 300 random draws of selection of training sets, and the final cross validation score is computed as the average of the likelihood of the test set. We also rank the model according to their Bayesian evidence, evaluated with the Laplace approximation (see Hara et al. 2020; Nelson et al. 2020).
In our analysis, we adopt as an eight-column matrix. Its columns are: an offset, the spectral shape indicators (i.e., stellar activity proxies): the SPIRou longitudinal magnetic field (), and the LBL-derived (Artigau et al. 2022, 2024) projection onto the 2nd derivative of the spectrum (d2v), projection onto the 3rd derivative of the spectrum (d3v), differential line width (dLW), and differential temperature with respect to a T=4000 K template (dTEMP4000) and, finally, a cosine and sine term at the sidereal day (period equal to 0.997 days). The grid of values chosen for the noise model is given in Table 19. All the possible combinations of the values of its components were generated, and the corresponding covariance matrices were created according to Eq. (4). In 31, we represent the histogram of cross validation scores.
The periodogram corresponding to the highest ranked model is represented in Fig. 11. It shows four clear peaks at 6, 3, 18 and 1.3 days. To evaluate the robustness of these detections to the noise model, for each the 20% highest ranked models, we report the false alarm probabilities of the peaks found. The false alarms of the highest ranking models are reported in Fig. 32. The percentage of cases where the 6, 3, 18 and 1.3 days signals were found in the 20% best models, the FAPs corresponding to the highest ranked model, and the median FAPs are reported in Table LABEL:tab:faps_cv. We find that the 6, 3, 18 days signal are present in a 100% of the highest ranked models, with median false alarm, probabilities of , and . The highest ranked models by Bayesian evidence give a similar picture, although they seem to provide more statistical significance for the signals (see Table 18). They can be considered respectively as a confirmed planet, a very strong candidate and a strong candidate.
@@
H.2 Apodized sine periodogram analysis.
@@
We compare the of a linear base model , and the of a model defined as the linear model of plus an apodized sinusoid . The ASP is defined as the difference between the two models
We use the same base model as in our analysis with the periodogram, and use the covariance which corresponds to the noise model with highest Bayesian evidence.
To compute the ASP, we consider a grid of timescales and , where is the total observation timespan of the data. In Fig. 34.a, left panel, we represent as defined in Eq. (5), maximised over for the values of in the grid. In other words, we always select the center of the time window which best fits the data.
We search iteratively for signals. Once the maximum peak of the ASP is found, then we include the corresponding model in the base linear model , and repeat the process. The four first iterations are shown in Fig. 34, a, b, c ,d respectively. The left panels represent the periodograms, the middle panels a zoom in on the highest peak, and the right hand panel represents a statistical significance test, explained below.
In Fig. 34.a, the dominating peak has a period of 6.02 days. We want to test whether the signal is statistically compatible with a constant one (i.e. with ). Denoting by the value of maximising the value of the periodogram (5) for a given and , we compute the distribution of
with the hypothesis that the model is correct, where the fitted cosine and sine amplitudes are obtained by fitting model to the data. can easily be expressed as a generalised distribution, with mean and variance is given by an analytical expression (Hara et al. 2022).
@@
Since the d3v vector, used in our linear base model , has shown in other SPIRou dataset to be less reliable202020https://lbl.exoplanets.ca/output-data we re-did the whole analysis with the and ASP periodograms without it. The only notable difference is that it seems to slightly decrease the stability of the 18.7 day signal in frequency, amplitude and phase. Let us note that the 2.99 day planet could be in 2:1 mean motion resonance with the 6 d planet. In Hara et al. (2019), it has been shown that eccentric planets can mimic 2:1 mean motion resonances. However, it is very unlikely the case here, as the 2.99 d signal is significant and sharp. In conclusion, both 18.7 and 2.99 d planets are viable candidates, but the evidence for the latter is stronger.

### v2

@@ Appendix C Gaussian process modeling of SPIRou’s longitudinal magnetic field
| GP of the SOPHIE optical RVs using SPIRou’s as an activity proxy | ||||
@@
Appendix D Corner plots and additional figures and tables of the one-planet model
Appendix E Corner plots and additional figures and tables of the RadVel two-planet model
@@
Appendix F Dynamical stability analysis
While somehow limited constraints on the orbital configuration can be obtained from the RV measurements, we performed a dynamical stability assessment for the two-planet system Gl 410b and 18.7-day candidate to consolidate the picture on the system architecture. The period ratio (larger than 3) is wide and Gl 410 b and the 18.7-day candidate are small enough that the system is safely stable for circular orbits and the nominal eccentricity values. Indeed the orbital separation in Hill radii is of which is far beyond the stability limit of (Gladman 1993). Although we favor a model with circular orbits, we can derive some constraints on the maximum stable eccentricities by computing the general Hill stability criterion as described in Petit et al. (2018). We take the RadVel posterior and compute the Angular Momentum Deficit (AMD) of the system as well as the critical value ensuring Hill stability (in other words, the maximum AMD such that close encounters are always forbidden). In Fig. 30, we compare the posterior eccentricity distribution resulting from the RadVel RV fit and the posterior eccentricity distribution only keeping the stable configurations. Rejecting Hill unstable configurations from the posterior, we obtain an upper limit for at 0.4 but no additional constraints on . Stable configurations are on average more circular. A similar analysis on the masses show that no stringer constraints can be drawn from the stability on this system.
The Hill stability can also constrain the maximum mutual inclination. To obtain this constraint, we assume that the system invariant plan is aligned with the line of sight252525As the main constraint comes from the mutual inclination and not the impact on the real planetary masses, this is a reasonable hypothesis. and we assign a random mutual inclination to each configuration in the posterior. Computing the critical AMD in this case, we find that systems with a mutual inclination smaller than 40 degrees are stable. As for the eccentricities, this is mainly an upper limit. The system is fully consistent with coplanar orbits.
Appendix G Additional radial velocity analysis of the SPIRou wPCA data
G.1 periodogram analysis
The periodogram has several inputs: the RV time series to be analyzed, its nominal uncertainties, a grid of frequencies , a linear model assumed to be in the data , an assumed covariance matrix of the noise as input. Overall, the model of the RV time series can be described by the equation
@@ Appendix G Additional radial velocity analysis of the SPIRou wPCA data
The periodogram looks for a small number of frequencies such that the model in Eq. (3) approximates the data. Depending on the different assumptions on the noise model and base linear model, the periodogram might find signals at different frequencies. To test the robustness of the results to a change in the model, we adopt the methodology of Hara et al. (2020). We consider a grid of potential noise models. For each of them, we computed the periodogram, and kept the signals that have a false alarm probability below 0.1. For each noise model in the grid, we then have candidate periodic signals. To each couple noise model/periodic signals, we attribute a cross-validation score. The model is fitted on 60% of the data points (the training set), and the likelihood of its prediction on the remaining 40% (the test set) is computed. This operation is repeated for 300 random draws of selection of training sets, and the final cross-validation score is computed as the average of the likelihood of the test set. We also rank the model according to their Bayesian evidence, evaluated with the Laplace approximation (see Hara et al. 2020; Nelson et al. 2020).
In our analysis, we adopt as an eight-column matrix. Its columns are: an offset, the spectral shape indicators (i.e., stellar activity proxies): the SPIRou longitudinal magnetic field (), and the LBL-derived (Artigau et al. 2022, 2024) projection onto the 2nd derivative of the spectrum (d2v), projection onto the 3rd derivative of the spectrum (d3v), differential line width (dLW), and differential temperature with respect to a T=4000 K template (dTEMP4000) and, finally, a cosine and sine term at the sidereal day (period equal to 0.997 days). The grid of values chosen for the noise model is given in Table 19. All the possible combinations of the values of its components were generated, and the corresponding covariance matrices were created according to Eq. (4). In 31, we represent the histogram of cross-validation scores.
The periodogram corresponding to the highest ranked model is represented in Fig. 11. It shows four clear peaks at 6, 3, 18 and 1.3 days. To evaluate the robustness of these detections to the noise model, for each the 20% highest ranked models, we report the false alarm probabilities of the peaks found. The false alarms of the highest ranking models are reported in Fig. 32. The percentage of cases where the 6, 3, 18 and 1.3 days signals were found in the 20% best models, the FAPs corresponding to the highest ranked model, and the median FAPs are reported in Table LABEL:tab:faps_cv. We find that the 6, 3, 18 days signal are present in a 100% of the highest ranked models, with median false alarm, probabilities of , and . The highest ranked models by Bayesian evidence give a similar picture, although they seem to provide more statistical significance for the signals (see Table 18). They can be considered respectively as a confirmed planet, a very strong candidate, and a strong candidate.
@@
G.2 Apodized sine periodogram analysis
@@
We compared the of a linear base model , and the of a model defined as the linear model of plus an apodized sinusoid . The ASP is defined as the difference between the two models
We used the same base model as in our analysis with the periodogram, and used the covariance that corresponds to the noise model with highest Bayesian evidence.
To compute the ASP, we consider a grid of timescales and , where is the total observation timespan of the data. In Fig. 34.a, left panel, we represent as defined in Eq. (5), maximized over for the values of in the grid. In other words, we always select the center of the time window which best fits the data.
We searched iteratively for signals. Once the maximum peak of the ASP was found, we then included the corresponding model in the base linear model , and repeated the process. The four first iterations are shown in Fig. 34, a, b, c, and d respectively. The left panels represent the periodograms, the middle panels a zoom in on the highest peak, and the right hand panel represents a statistical significance test, explained below.
In Fig. 34.a, the dominating peak has a period of 6.02 days. We want to test whether the signal is statistically compatible with a constant one (i.e., with ). Denoting by the value of maximizing the value of the periodogram (5) for a given and , we compute the distribution of
with the hypothesis that the model is correct, where the fitted cosine and sine amplitudes are obtained by fitting model to the data. can easily be expressed as a generalized distribution, with mean and variance is given by an analytical expression (Hara et al. 2022).
@@
Since the d3v vector, used in our linear base model , has shown in other SPIRou data set to be less reliable262626https://lbl.exoplanets.ca/output-data we re-did the whole analysis with the and ASP periodograms without it. The only notable difference is that it seems to slightly decrease the stability of the 18.7 day signal in frequency, amplitude and phase. We note that the 2.99 day planet could be in 2:1 mean motion resonance with the 6 d planet. In Hara et al. (2019), it has been shown that eccentric planets can mimic 2:1 mean motion resonances. However, it is very unlikely the case here, as the 2.99 d signal is significant and sharp. In conclusion, both 18.7 and 2.99 d planets are viable candidates, but the evidence for the latter is stronger.